Two cylindrical rods, AC made of aluminum and CD made of steel, are joined at Cand restrained by rigid supports at A and D. I1.Determine the reactions at A and D. (in kips)

2. The deflection of point C.

Answer all parts of the question please
Answer #1

Below i have attached solutions for this

AB RA o foafele ¡ RD DAB+ DB C + D CD =0 deflection , in , enca én codA B b con AAB= YRA (a) rda ta DA BC= 4 (RA -34)(6) rida

Ra (o-8) +(RA - 34) = - ( Ro-18) (9/8) (1318)2 (29/1014) 18 Ra -34 = - Ra -4 800- (1.8 +0.1727RA = 34 +48 X0:17 2 RA= 21,43 k